OpenStudy (anonymous):

Half the supplement of an angle is 18 degrees less than four times the complement of the angle. Set up and solve an equation to find the measure of the angle, the complement, and supplement.

OpenStudy (mertsj):

Let x = the angle then 180-x= supplement and 90-x = complement.

OpenStudy (anonymous):

so far the equation I have is 1/2(18)-x=4(90-x)+18 is that right?

OpenStudy (anonymous):

@mertsj

OpenStudy (anonymous):

180 and 18

OpenStudy (mertsj):

\[\frac{1}{2}(180-x)=4(90-x)-18\]

OpenStudy (anonymous):

x=84 ?

OpenStudy (anonymous):

@mertsj

OpenStudy (yttrium):

x=72

OpenStudy (anonymous):

how did you get that?

OpenStudy (yttrium):

Just tell us how you've got 84

OpenStudy (anonymous):

1/2(180)-x=4(90-x)-18 -->> 90-x=360-4x-18 +4x +4x 90+3x=342 -90 -90 3x=252 x=84

OpenStudy (anonymous):

where do i go wrong

OpenStudy (yttrium):

You forgot to divide the x to 2 in the left side of the eq'n. It should be 90- (x/2) = 4(90-x)- 18 But I have a simpler solution, maybe you can solve that easier.

OpenStudy (anonymous):

i dont understand exactly so it would be 90-(x/2)=342-4x

OpenStudy (anonymous):

wait ok i get it now

OpenStudy (anonymous):

so complement is 18 supplement is 108

OpenStudy (yttrium):

it says here that the half of the supplement, which is 1/2 (180-x) is equal to 18 deg less than four times of its complement which in symbols will be equal to 4(90-x)-18 You get it?

OpenStudy (anonymous):

54

OpenStudy (yttrium):

How you come up to 54? Do you want me to show my solution?

OpenStudy (anonymous):

yes cause i cant get it

OpenStudy (yttrium):

\[(1/2) (180-x) = 4(90-x) - 18\] \[(1/2) (180-x) = 360 - 4x - 18\] \[(1/2) (180 - x) = 342 - 4x\] \[180-x = 684 - 8x\] Therefore, \[ 7x = 504\] x = 72

OpenStudy (anonymous):

yes i got that part which is the angle, but i also need to find the supplement and complement but for that do i just replace x with 72 thats how i got 54

OpenStudy (yttrium):

Well, it's still not 54. The complement is just 90-x and the supplement is just 180-x

OpenStudy (anonymous):

18 & 108

OpenStudy (yttrium):

Nice. You got it right now! :D

OpenStudy (anonymous):

thanks, im bad at math so it takes awhile. but thank you for all your help